OpenStudy (anonymous):

factor using foil method. (9+8i)(9-8i)

OpenStudy (anonymous):

64i^2+142i+81

OpenStudy (anonymous):

hold on not right

OpenStudy (anonymous):

Easy .First 9*9 + 9*(-8i) + 8i*9 + 8i*(-8i) Change i square to -1 This is answer 81 -72i + 72i + 64 = 145

OpenStudy (anonymous):

first, is i imaginary?

OpenStudy (anonymous):

or is it just a number?

OpenStudy (anonymous):

can anyone try to help me with mine??

OpenStudy (anonymous):

You also do that , but can check in (a + bi) (c+di), if a*d =-c*d and b*d=-X, This mean the equation is a numeric

OpenStudy (anonymous):

I can